OpenStudy (lgbasallote):

\(\huge \mathtt{\text{lalaly users only!}}\) \(\mathsf{\text{(for now :P)}}\) integrate using partial fractions \(\LARGE \int \frac{e^x}{(e^x - 1)(e^x + 3)} dx\)

OpenStudy (lgbasallote):

@lalaly

OpenStudy (lalaly):

lol sorry it took me a long time,,, i was solving another integral ...

OpenStudy (lgbasallote):

nahh izfine that one was difficult :p haha

OpenStudy (lgbasallote):

i made this post lalaly users only :p you should be flattered -__- haha jk

OpenStudy (dumbcow):

users ? are there more than 1 :)

OpenStudy (lgbasallote):

exactly =))

OpenStudy (lgbasallote):

@dumbcow why'd you reply here...now @lalaly is not going to answer my question :C

OpenStudy (lalaly):

give me 5 more min the page crashed and lol i will do it dont worry,,, dumbcow is always welcome xD

OpenStudy (lgbasallote):

oh..i thought you wont reply anymore coz you stopped typing :P hahaha

OpenStudy (dumbcow):

haha..ok i will delete nevermind...thanks lalaly

OpenStudy (lalaly):

let u=e^x du=e^xdx so the integral becomes\[\int\limits{\frac{du}{(u-1)(u+3)}}\] first thing u do is break \[\frac{1}{(u-1)(u+3)}\]into partial fractions \[\frac{1}{(u-1)(u+3)}=\frac{A}{u-1}+\frac{B}{u+3}\] now you multiply both sides by (u-1)(u+3) \[\cancel{(u-1)(u+3)}\times \frac{1}{\cancel{(u-1)(u+3)}}=(\frac{A}{u-1}+\frac{B}{u+3})\times(u-1)(u+3)\] now we have 1=A(u+3)+B(u-1) now to find A and B from (u-1)(u+3) u tkae first u=1 and substitute it in the equation 1=A(1+3)+B(1-1) 1=4A+0 so A=1/4 then u take u=-3 and u do the asme 1=A(-3+3)+B(-3-1) 1=0-4B so B=-1/4 now that u know A and B \[\frac{1}{(u-1)(u+3)}=\frac{\frac{1}{4}}{u-1}+\frac{\frac{-1}{4}}{u+3}\] \[\frac{1}{(u-1)(u+3)}=\frac{1}{4(u-1)}-\frac{1}{4(u+3)}\]

OpenStudy (lalaly):

now substitute that in the integral \[\int\limits{\frac{du}{(u-1)(u+3)}}=\int\limits{(\frac{1}{4(u-1)}-\frac{1}{4(u+3)})du}\]

OpenStudy (lalaly):

\[=\int\limits{\frac{du}{4(u-1)}}-\int\limits{\frac{du}{4(u+3)}}\]

OpenStudy (lalaly):

now its an easy integral :D

OpenStudy (lalaly):

tell me if there are things u dont understand:D

OpenStudy (lalaly):

one more thing after u find the integral u substitute u=e^x back xD

OpenStudy (lalaly):

do it and i will tell u if its wrong or right

OpenStudy (lgbasallote):

\(\large \frac{\ln (u-1)}{4} - \frac{\ln (u+3)}{4}\) \(\Large \frac{\ln \frac{(u-1)}{(u+3)}}{4}\) \(\LARGE \ln \sqrt[4]{\frac{(u-1)}{(u+3)}}\) \(\LARGE \ln \sqrt[4]{\frac{(e^x - 1)}{(e^x + 3)}}\)

OpenStudy (lalaly):

yeah +C xD

OpenStudy (lgbasallote):

oh yeah that too :P

OpenStudy (lgbasallote):

thanks :D that was pretty hard gotta practice more! then ill learn partial fractions + trig sub

OpenStudy (lalaly):

cool im here if u need any more help

OpenStudy (lgbasallote):

well...now that you mention it...

OpenStudy (lgbasallote):

can you do this: \(\LARGE \int \frac{x^4 - x^2 + 9}{x(x^2 + 3x + 3)^2} dx\) looks hard :/ is it?

OpenStudy (lalaly):

lol same way i showd first thing factorise the denominator

OpenStudy (lgbasallote):

can i try? :DD

OpenStudy (lgbasallote):

\(\LARGE \frac{A}{x} + \frac{Bx + C}{x^2 -3x + 3} + \frac{Dx + E}{(x^2 - 3x + 3)^2} \) that right?

OpenStudy (lgbasallote):

the middle terms are +3x lol

OpenStudy (lalaly):

yeah do the rest

OpenStudy (lgbasallote):

A(x^2 + 3x + 3) + Bx + C(x^3 + 3x^2 + 3x) + Dx^2 + Ex = x^4 - x^2 + 9 Ax^2 + 3Ax + 3A + Bx^4 + 3Bx^3 + 3Bx^2 + Cx^3 + 3Cx^2 + 3Cx + Dx^2 + Ex = x^4 - x^2 + 9

OpenStudy (lgbasallote):

3A = 9 A = 3

OpenStudy (lgbasallote):

3A + 3C = 0 9 + 3C = 0 3C = -9 C = -3

OpenStudy (lgbasallote):

coefficient of x^3 3B+ C = 0 3B - 3 = 0 3B = 3 B = 1

OpenStudy (lgbasallote):

the previous ones were the constant and the coefficient of x respectively

OpenStudy (lgbasallote):

coefficient of x^2 A + 3B + 3C + D = -1 3 + 3 - 9 + D = -1 D = -1 + 9 - 3 -3 D = 2

OpenStudy (lgbasallote):

oh no! i forgot Exx im ruined T_T

OpenStudy (lalaly):

lol solve it all and we'll see where u go wrong

OpenStudy (lgbasallote):

okay :(